find the width of a rectangular garden if the perimeter is 66 feet and the length is 7 feet longer than the width.
Answer by ikleyn(52817)

.

The perimeter = W + (W+7) + W + (W+7) = 66


2W + (2W + 14) = 66


4W = 66 - 14 = 52


W = 52/4 = 13 feet.      ANSWER

Solved.